Status Report

NOAA SATOPS Morning Report: Friday, August 11, 2006

By SpaceRef Editor
August 11, 2006
Filed under , , ,

JDay 223

GOES

All GOES operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For GOES status and other information please see: http://www.oso.noaa.gov/goes/index.htm

All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

POES

N14 J223 14:41Z – 20:14Z Successfully performed attitude maneuver to point MSU instrument into deep space for one orbit in support of ORA instrument calibration activity.

NOTE to N14 AVHR & HIRS users: Three of the AVHR detectors and the HIRS electronics were turned off following the maneuver to allow for thermal rebalancing. AVHR detectors and the HIRS electronics are scheduled to be turned back on Monday morning.

All POES operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For a detailed status and description of POES see: http://www.oso.noaa.gov/poes/index.htm

DMSP

All DMSP operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For more DMSP Information please see: http://dmsp.ngdc.noaa.gov/dmsp.html

Questions or comments to Keith.Amburgey@noaa.gov

SpaceRef staff editor.